# LeanQuant: Accurate and Scalable Large Language Model Quantization with Loss-error-aware Grid

### v2

@@ Abstract
Large language models (LLMs) have shown immense potential across various domains, but their high memory requirements and inference costs remain critical challenges for deployment. Post-training quantization (PTQ) has emerged as a promising technique to reduce memory requirements and decoding latency. However, recent accurate quantization methods often depend on specialized computations or custom data formats to achieve better model quality, which limits their compatibility with popular frameworks, as they require dedicated inference kernels tailored to specific hardware and software platforms, hindering wider adoption. Furthermore, many competitive methods have high resource requirements and computational overhead, making it challenging to scale them to hundreds of billions of parameters. In response to these challenges, we propose LeanQuant (Loss-error-aware Network Quantization), a novel quantization method that is accurate, versatile, and scalable. In the existing popular iterative loss-error-based quantization framework, we identify a critical limitation in prior methods: the min-max affine quantization grid fails to preserve model quality due to outliers in inverse Hessian diagonals. To overcome this fundamental issue, we propose learning loss-error-aware grids, instead of using non-adaptive min-max affine grids. Our approach not only produces quantized models that are more accurate but also generalizes to a wider range of quantization types, including affine and non-uniform quantization, enhancing compatibility with more frameworks. Extensive empirical evaluations on recent LLMs demonstrate that LeanQuant is highly accurate, comparing favorably against recent competitive baselines in model quality, and scalable, achieving very accurate quantization of Llama-3.1 405B, one of the largest open-source LLMs to date, using two Quadro RTX 8000-48GB GPUs in 21 hours.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive reasoning (Wei et al., 2022) and problem solving abilities (Kojima et al., 2022), and have shown the potential to bring transformative changes to various fields such as law (Kaddour et al., 2023), education (Kasneci et al., 2023), and medicine (Thirunavukarasu et al., 2023). However, deploying LLMs in a cost-effective manner presents significant challenges due to their substantial memory and computational demands (Chen et al., 2023), which hinders the accessibility and democratization of artificial intelligence (AI) (Kaddour et al., 2023).
@@
Post-training quantization (PTQ) (Krishnamoorthi, 2018) is a technique for reducing the memory requirement for model inference by reducing the precision of floating-point weights of a pre-trained model and storing them in a compact low-bit-width format. PTQ offers the additional benefit of reducing the decoding latency of LLMs by reducing memory reads, since LLM inference is often bottlenecked by memory bandwidth (Kim et al., 2023). Although quantization causes a certain amount of precision loss in the parameters, the model quality can be reasonably preserved even in lower bit widths (Frantar et al., 2022; Chee et al., 2024). For many tasks, a quantized model is preferred over a full model due to its better size-accuracy trade-off (Dettmers & Zettlemoyer, 2023). As open-source foundational models continue to scale up in size (Dubey et al., 2024), accurate and efficient quantization becomes essential for making AI accessible to a wider audience. For instance, serving Llama-3.1 405B (Dubey et al., 2024) using its original 16-bit weights requires a cluster of 2 nodes, each with 8$\times$80GB GPUs, while the 4-bit quantized version can be deployed on a single nodes of 8$\times$48GB GPUs, eliminating the overhead of inter-node communication.
Challenges of Deploying Quantized Models One of the biggest challenges of successful deployment of quantized models is implementing optimized kernels for quantized GEMM (general matrix multiply) that are tailored to various hardware platforms and software frameworks. In order to accelerate inference of quantized models, fused kernels, which fuse dequantization and matrix multiplication in the same subroutine, have to be implemented and tuned for the specific hardware accelerator. These kernels require specialized designs and tunings for different hardware accelerators to be fully optimized (Park et al., 2022).
@@
Recent quantization algorithms have chosen to employ specialized computations or custom data formats to reduce the impact of quantization on model quality, but they require more sophisticated kernel designs for efficient inference. For example, AQLM (Egiazarian et al., 2024) and QUIP# (Tseng et al., 2024) perform dequantization through look-ups from multi-dimensional or multi-bit codebooks, and Dotzel et al. (2024) proposed new data types such as Student Float to reduce quantization errors. While these approaches demonstrate promising results, their reliance on specialized operations and data formats can hinder their widespread adoption due to the need for optimized inference kernels for each hardware platform and software framework. For example, llama.cpp (Gerganov, 2023), a popular LLM inference engine that supports mobile devices, only supports affine and non-uniform quantization formats. Consequently, instead of focusing on developing better quantization methods with specialized operations, it may be more worthwhile to investigate improving the accuracy of existing widely adopted quantization formats, such as affine integer quantization and non-uniform quantization, which are supported by popular deep learning libraries (Paszke et al., 2019) and deployment frameworks (Kwon et al., 2023).
@@
Scalability Challenges of Accurate Quantization To improve the quality of quantized models, existing approaches often incur higher computational overhead and require more hardware resources. As foundational models scale up in size (Hoffmann et al., 2022), these quantization approaches may struggle to scale to very large models such as Llama-3.1 405B (405 billion parameters) (Dubey et al., 2024). For instance, LLM-QAT (Liu et al., 2023) uses 100K samples of training data and hundreds of GPU-hours to recover the performance of a quantized LLaMA-13B model (Touvron et al., 2023a). For AQLM (Egiazarian et al., 2024), the time needed for quantizing a 7B to 70B LLM ranges from 1 to 14 days of an A100-80GB GPU. For SqueezeLLM (Kim et al., 2023), due to its use of the gradients of model parameters, quantizing a 70B LLM requires at least 240GB of total GPU memory, or 8$\times$32GB GPUs. As these accurate quantization approaches demand significant hardware resources and long optimization time, it is crucial to develop accurate methods that are efficient in terms of resource usage and time cost, to ensure the accessibility of increasingly large foundational models.
@@
Our Proposal In this work, we propose LeanQuant, an accurate, versatile, and scalable quantization approach. We build upon the iterative loss-error-based quantization framework (Frantar & Alistarh, 2022; Frantar et al., 2022) and identify one of the biggest limitations of such methods: the min-max affine quantization grid introduces high loss errors due to the existence of outliers in the inverse Hessian diagonals. We introduce techniques for learning loss-error-aware quantization grids, which mitigate this issue and greatly improve the accuracy and quality of quantized models. We empirically demonstrate that LeanQuant compares favorably against competitive baselines in the 4/3/2-bit regions. Our approach is versatile, able to generalize to multiple commonly used quantization formats, such as affine and non-uniform quantization, allowing our quantized models to be directly compatible with existing highly optimized inference kernels (Frantar et al., 2024; Park et al., 2022) for maximum accessibility. Furthermore, our method is scalable and efficient. By designing and implementing a fused GPU kernel for LeanQuant grid learning, we achieve the accurate quantization of LLMs up to 123B in size using a single L40s-48GB GPU in 4 hours, and Llama-3.1 405B using 2 Quadro RTX 8000-48GB GPUs in 21 hours.
@@ 2.1 Quantization Grid
Quantization achieves model compression by representing the full-precision floating-point parameters with a limited set of grid points on the quantization grid. The number of points on the grid depends on the bit width: a $b$-bit code represents $2^{b}$ distinct values, so 2-bit quantization yields 4 grid points. The placement of these points significantly affects the precision of the original parameters, and imprecise placement can degrade model quality. To address this, different types of quantization grids, such as affine, non-uniform, and normal, have been introduced, which we survey as follows.
Affine Grid In an affine quantization grid (Krishnamoorthi, 2018), the grid points are evenly spaced over the dynamic range of a group of weights. The weights of the network are divided into groups of fixed size, e.g. every 128 contiguous parameters as a group. For min-max asymmetric affine quantization, the scaling factor $S$ and the zero-point $Z$ are recorded for each group of weights (the zero-point is omitted in the symmetric affine case). Then, the $i$-th weight $w_{i}$ in a group $\mathbf{w}$ is quantized to $b$-bit as follows,
@@
[…]
Non-uniform Grid The grid points on a non-uniform grid are placed in a non-equidistant manner (Li et al., 2019). The motivation behind non-uniform quantization is to allow for finer precision in regions where model parameters are more concentrated or sensitive. Each row in a weight matrix has a distinct set of grid points $\mathcal{G}$, where $\lvert\mathcal{G}\rvert=2^{b}$ for $b$-bit non-uniform quantization. The weight $w_{i}$ is quantized to the nearest grid point in $\mathcal{G}$ as follows,
@@
Other Grid Types Previous works have observed that LLM parameters tend to be distributed similarly as Normal or Student T’s Distribution, hence they propose new grid types, such as NormalFloat (Dettmers et al., 2024) and Student Float (Dotzel et al., 2024), which place grid lines at the quantiles of these distributions. Our proposed method can also be extended to work with these quantization formats.
@@ 2.2 Iterative Loss-error-based Quantization
Iterative loss-error-based quantization (Frantar & Alistarh, 2022) is a promising framework for quantizing deep neural networks to low bit widths while maintaining strong performance on downstream tasks. In particular, Optimal Brain Quantization (OBQ) (Frantar & Alistarh, 2022), which is based on the seminal works by LeCun et al. (1989) and Hassibi et al. (1993), aims to minimize the impact of weight perturbations introduced by parameter quantization on the network’s task loss.
Let $\mathcal{L}(\mathbf{w}_{\mathcal{N}})$ be the task loss of a network $\mathcal{N}$ evaluated at the weights $\mathbf{w}_{\mathcal{N}}$ (flattened to a vector). Then, the OBQ objective is to minimize the loss error $\epsilon$, which is defined as
@@
where $\bm{\delta}_{\mathcal{N}}$ is the weight perturbation introduced by quantization. The loss error $\epsilon$ can be approximated with a Taylor series (LeCun et al., 1989) as
@@
where the first term is omitted due to $\frac{\partial{\mathcal{L}}}{\partial{\mathbf{w}_{\mathcal{N}}}}\approx\mathbf{0}$ in a converged network, and the third and higher terms can be ignored due to small norms. Computing the exact Hessian $\mathbf{H}=\frac{\partial^{2}{\mathcal{L}}}{\partial{\mathbf{w}^{2}_{\mathcal{N}}}}$ in a deep network is difficult, hence OBQ leverages an approximation of loss error proposed by Nagel et al. (2020),
@@
where the Hessian $\mathbf{H}=2\mathbf{X}\mathbf{X}^{\top}$ (from Equation 1) is computed on a small sample of input data and $e_{i}$ is the $i$-th standard basis vector. Solving Equation 2 yields the optimal weight perturbation $\bm{\delta}_{i}$ and the resulting loss error $\epsilon_{i}$ after quantizing $w_{i}$,
@@
The loss error $\epsilon_{i}$ reflects the negative impact of quantizing parameter $w_{i}$ on the model quality, and it is always a non-negative value. OBQ utilizes the loss error $\epsilon_{i}$ as a heuristic for greedy optimization. Specifically, in each iteration, OBQ computes $\epsilon$ for all weights in a row and greedily selects the $i$-th parameter with the minimum $\epsilon_{i}$ to quantize. Then, $w_{i}$ is round to the nearest value on the quantization grid, and the remaining weights are updated via $\mathbf{w}\leftarrow\mathbf{w}-\bm{\delta}_{i}$, and the updated inverse Hessian for the remaining weights, with the $i$-th row and column removed from $\mathbf{H}$, is computed as
This iterative process continues until all weights are quantized.
Scaling to Billion-Parameter LLMs Using Cholesky and Dampening OBQ produces accurate post-training quantized models for million-parameter networks, but fails to scale to billion-parameter LLMs due to two primary reasons: the inefficient time complexity and the accumulation of numerical inaccuracies during updates. To improve its computational efficiency, Frantar et al. (2022) propose to quantize the weights in a fixed non-greedy order for all rows, and keep the weight updates within a block of $B$ columns at a time. To prevent model quality collapse resulted from the accumulation of numerical inaccuracies by repeated weight updates, Frantar et al. (2022) propose to apply a mild dampening (1% of the average diagonals) to the diagonal entries of the Hessian $\mathbf{H}$ and leverage a Cholesky decomposition of the inverse Hessian $\mathbf{H}^{-1}$ in place of the update in Equation 4. The resulting algorithm is GPTQ, which is able to efficiently quantizes billion-parameter LLMs.
@@ 3 Methodology
Figure 1: (Left) The empirical distributions of inverse Hessian diagonals, computed on 262K tokens from the C4 dataset for the Llama-3-8B model, contain outliers that can cause high loss errors. (Right) Our proposed loss-error-aware non-uniform and affine grids better preserves the quantized precision of outliers, leading to more accurate quantized models.
@@ 3.1 Revisiting the Loss Error
To motivate our proposed approach, we first revisit the loss error $\epsilon_{i}$ in Equation 3, which approximates the (detrimental) increase in the network’s task loss, introduced by quantizing weight $w_{i}$. This error $\epsilon_{i}$ has been used as a heuristic in multiple previous works (LeCun et al., 1989; Hassibi et al., 1993; Singh & Alistarh, 2020; Frantar & Alistarh, 2022) for choosing the next best weight $i$ to prune or quantize. It has been shown to be a highly informative metric for measuring the impact of quantization.
@@
Hence, we further examine the empirical distribution of $\frac{1}{\mathrm{diag}(\mathbf{H}^{-1})}$, to which the loss error row vector $\bm{\epsilon}$ is proportional. We obtain the empirical distributions on layers of Llama-3-8B (Dubey et al., 2024) with 128 sequences of length 2048 tokens from the C4 dataset (Raffel et al., 2020), and compute the inverse Hessian as $\mathbf{H}^{-1}=(2\mathbf{X}\mathbf{X}^{\top})^{-1}$ where $\mathbf{X}$ is the layer input matrix. As shown in Figure 1, The majority of the inverse diagonals are concentrated in low-magnitude regions, with a few outliers having high magnitudes. Quantizing the weights corresponding to these outliers can lead to high loss errors if these weights are not well-aligned with the quantization grid points. Preserving the quantized precision of the weights corresponding to these inverse-diagonal outliers is especially important because the loss error increases quadratically with their quantization error (Equation 5). Iterative loss-error-based quantization approaches (OBQ, GPTQ, etc.) employ min-max affine quantization grid, which is suboptimal for preserving the quantized precision of the inverse-diagonal outliers, leading to high loss errors and model quality degradation. Our idea is to learn quantization grids that minimize the loss error $\epsilon$.

### v3

@@ Abstract
Large language models (LLMs) have shown immense potential across various domains, but their high memory requirements and inference costs remain critical challenges for deployment. Post-training quantization (PTQ) has emerged as a promising technique to reduce memory requirements and decoding latency. However, recent accurate quantization methods often depend on specialized computations or custom data formats to achieve better model quality, which limits their compatibility with popular frameworks, as they require dedicated inference kernels tailored to specific hardware and software platforms, hindering wider adoption. Furthermore, many competitive methods have high resource requirements and computational overhead for quantizing models, making it challenging to scale them to hundreds of billions of parameters. In response to these challenges, we propose LeanQuant (Loss-error-aware network Quantization), a novel quantization method that is accurate, versatile, and scalable. In the existing popular iterative loss-error-based quantization framework, we identify a critical limitation in prior methods: the min-max affine quantization grid fails to preserve model quality due to outliers in inverse Hessian diagonals. To overcome this fundamental issue, we propose learning loss-error-aware grids, instead of using non-adaptive min-max affine grids. Our approach not only produces quantized models that are more accurate but also generalizes to a wider range of quantization types, including affine and non-uniform quantization, enhancing compatibility with more frameworks. Extensive experiments with recent LLMs demonstrate that LeanQuant is highly accurate, comparing favorably against competitive baselines in model quality, and scalable, achieving very accurate quantization of Llama-3.1 405B, one of the largest open-source LLMs to date, using two Quadro RTX 8000-48GB GPUs in 21 hours. Our code is available at https://github.com/LeanModels/LeanQuant.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive reasoning (reasoning) and problem solving abilities (zero_shot), and have shown the potential to bring transformative changes to various fields such as law (app_llm), education (llm_education), and medicine (llm_medicine). However, deploying LLMs in a cost-effective manner presents significant challenges due to their substantial memory and computational demands (frugalgpt), which hinders the accessibility and democratization of artificial intelligence (AI) (app_llm).
@@
Post-training quantization (PTQ) (uniform_quant) is a promising technique for reducing the memory footprint of model inference by lowering the precision of a pre-trained model’s parameters and storing them in a compact, low-bit-width format. PTQ offers the additional benefit of reducing the decoding latency of LLMs by reducing memory reads, since LLM inference is often bottlenecked by memory bandwidth (sqllm). Although quantization causes a certain amount of precision loss in the parameters, the model quality can be reasonably preserved even in lower bit widths (gptq; quip). For many tasks, a quantized model is preferred over a full model due to its better size-accuracy trade-off (scaling_law_4bit). As open-source foundational models continue to scale up in size (llama3), accurate and efficient quantization becomes essential for making AI accessible to a wider audience. For instance, serving Llama-3.1 405B (llama3) with its original 16-bit weights requires a cluster of two nodes, each equipped with 8×80GB GPUs. In contrast, a 4-bit quantized version can be deployed on a single node with 8×48GB GPUs, eliminating inter-node communication overhead.
@@
Challenges of Deploying Quantized Models One of the biggest challenges of successful deployment of quantized models is implementing optimized kernels for quantized GEMM (general matrix multiply) that are tailored to various hardware platforms and software frameworks. In order to accelerate inference of quantized models, fused kernels, which fuse dequantization and matrix multiplication in the same subroutine, have to be implemented and tuned for the specific hardware accelerator. These kernels require specialized designs and tunings for different hardware accelerators to be fully optimized (lut_gemm). Recent quantization algorithms have chosen to employ specialized computations or custom data formats to reduce the impact of quantization on model quality, but they require more sophisticated kernel designs for efficient inference. For example, AQLM (aqlm) and QUIP# (quipsharp) perform dequantization through look-ups from multi-dimensional or multi-bit codebooks, and student_t proposed new data types such as Student Float to reduce quantization errors. While these approaches demonstrate promising results, their reliance on specialized operations and data formats can hinder their widespread adoption due to the need for optimized inference kernels for each hardware platform and software framework. For example, llama.cpp (llamacpp), a popular LLM inference engine that supports mobile devices, only supports affine and non-uniform quantization formats. Consequently, instead of focusing on developing better quantization methods with specialized operations, it may be more worthwhile to investigate improving the accuracy of existing widely adopted quantization formats, such as affine integer quantization and non-uniform quantization, which are supported by popular deep learning libraries (pytorch) and deployment frameworks (vllm).
@@
Scalability Challenges of Accurate Quantization To improve the quality of quantized models, existing approaches often incur higher computational overhead and require more hardware resources. As foundational models scale up in size (scaling_law), these quantization approaches may struggle to scale to very large models, such as Llama-3.1 405B (405 billion parameters) (llama3). For instance, LLM-QAT (llm_qat) uses 100K samples of training data and hundreds of GPU-hours to recover the performance of a quantized LLaMA-13B model (llama). For AQLM (aqlm), the time needed for quantizing a 7B to 70B LLM ranges from 1 to 14 days of an A100-80GB GPU. For SqueezeLLM (sqllm), due to its use of the gradients of model parameters, quantizing a 70B LLM requires at least 240GB of total GPU memory. Given the significant hardware resources and lengthy optimization times of these quantization approaches, developing accurate yet efficient methods is crucial for ensuring accessibility of larger foundational models.
@@
Our Proposal In this work, we propose LeanQuant, an accurate, versatile, and scalable quantization approach. We build upon the iterative loss-error-based quantization framework (obq; gptq) and identify one of the biggest limitations of such methods: the min-max affine quantization grid introduces high loss errors due to the existence of outliers in the inverse Hessian diagonals. We introduce techniques for learning loss-error-aware quantization grids, which mitigate this issue and greatly improve the quality of quantized models. We empirically demonstrate that LeanQuant compares favorably against competitive baselines in the 4/3/2-bit regions. Our approach is versatile, able to generalize to multiple commonly used quantization formats, such as affine and non-uniform quantization, allowing our quantized models to be directly compatible with existing highly optimized inference kernels (marlin; lut_gemm) for maximum accessibility. Furthermore, our method is scalable and efficient. By designing and implementing a fused GPU kernel for LeanQuant grid learning, we achieve the accurate quantization of LLMs up to 123B in size using a single L40s-48GB GPU in 4 hours, and Llama-3.1 405B using 2 Quadro RTX 8000-48GB GPUs in 21 hours.
@@ 2.1 Quantization Grid
Quantization enables model compression by representing full-precision floating-point parameters with a limited set of grid points on a quantization grid. The number of available grid points is determined by the bit width: a $b$-bit code allows for $2^{b}$ distinct values, meaning 2-bit quantization results in four grid points. A visual explanation of quantization grid can be found in Appendix A. The placement of these points is crucial, as inaccurate placements can degrade model quality. To mitigate this, various quantization grids—such as affine, non-uniform, and normal—have been proposed. We provide an overview of these approaches below.
Affine Grid In an affine quantization grid (uniform_quant), the grid points are evenly spaced between the minimum and maximum of a set of weights. To achieve finer quantization precision, the network’s weights are divided into groups (e.g., every 128 contiguous parameters). In min-max asymmetric affine quantization, each weight group is associated with a scaling factor $S$ and a zero-point $Z$ (with $Z$ omitted in the symmetric case). The $i$-th weight $w_{i}$ in a group $\mathbf{w}$ is quantized to a $b$-bit integer $w_{i}^{\mathbf{int}}$ as follows:
@@
where $\lfloor\cdot\rceil$ denotes rounding, $\mathrm{clip}(\cdot)$ ensures the value remains within the $b$-bit integer range, and $\mathrm{quant}_{\textit{aff}}(w_{i},S,Z)$ is the quantized value of $w_{i}$.
Non-uniform Grid The grid points on a non-uniform grid are placed in a non-equidistant manner (nonuniform). The motivation behind non-uniform quantization is to allow for finer precision in regions where model parameters are more concentrated or sensitive. Each row in a weight matrix has a distinct set of non-uniform grid points $\mathcal{G}$, where $\lvert\mathcal{G}\rvert=2^{b}$ for $b$-bit quantization. The weight $w_{i}$ is quantized to the nearest grid point in $\mathcal{G}$ as follows,
@@
Other Grid Types Previous works have observed that the distribution of LLM parameters often resembles Normal or Student T’s Distribution. Consequently, grid types such as NormalFloat (qlora) and Student Float (student_t) have been proposed, which align grid points with quantiles of these distributions. Our proposed method can be extended to support them.
@@ 2.2 Iterative Loss-error-based Quantization
Iterative loss-error-based quantization (obq) is a promising framework for quantizing deep neural networks to low bit widths while preserving model quality. In particular, Optimal Brain Quantization (OBQ) (obq), which is based on the seminal works by obd and obs, aims to minimize the impact of weight perturbations introduced by parameter quantization on the network’s task loss. Let $\mathcal{L}(\mathbf{w}_{\mathcal{N}})$ be the task loss of a network $\mathcal{N}$ evaluated at its weights $\mathbf{w}_{\mathcal{N}}$ (flattened to a vector). Then, the OBQ objective is to minimize the loss error $\epsilon$, which is defined as
@@
where $\bm{\delta}_{\mathcal{N}}$ is the weight perturbation introduced by quantization. The loss error $\epsilon$ can be approximated with a Taylor series (obd) as
@@
where the first term is omitted due to $\frac{\partial{\mathcal{L}}}{\partial{\mathbf{w}_{\mathcal{N}}}}\approx\mathbf{0}$ in a converged network, and the third and higher terms can be ignored due to small norms. Computing the exact Hessian $\mathbf{H}=\frac{\partial^{2}{\mathcal{L}}}{\partial{\mathbf{w}^{2}_{\mathcal{N}}}}$ in a deep network is difficult, hence OBQ leverages an approximation of loss error proposed by adaround,
@@
where $e_{i}$ is the $i$-th standard basis vector and $\mathbf{H}=2\mathbf{X}\mathbf{X}^{\top}$ is the Hessian from Equation 1 (computed on a small sample of input data). Solving Equation 2 yields the optimal weight perturbation $\bm{\delta}_{i}$ and loss error $\epsilon_{i}$ after quantizing $w_{i}$,
@@
The loss error $\epsilon_{i}$ quantifies the degradation in model quality caused by quantizing parameter $w_{i}$ and is always non-negative. OBQ leverages $\epsilon_{i}$ as a heuristic for greedy optimization. In each iteration, OBQ computes $\epsilon$ for all weights in a row and greedily selects the parameter $w_{i}$ with the smallest $\epsilon_{i}$ for quantization. The selected parameter is then rounded to the nearest value on the quantization grid, and the remaining weights are updated as $\mathbf{w}\leftarrow\mathbf{w}-\bm{\delta}_{i}$. This iterative process continues until all weights are quantized.
Scaling to Billion-Parameter LLMs Using Cholesky and Dampening OBQ produces accurate post-training quantized models for million-parameter networks, but fails to scale to billion-parameter LLMs due to two primary reasons: the inefficient time complexity and the accumulation of numerical inaccuracies during updates. To improve its computational efficiency, gptq propose to quantize the weights in a fixed non-greedy order for all rows, and keep the weight updates within a block of $B$ columns at a time. To prevent model quality collapse from the accumulation of numerical inaccuracies by repeated weight updates, gptq propose to apply a mild dampening (1% of the average diagonals) to the diagonal Hessian and leverage a Cholesky decomposition to compute the inverse Hessian $\mathbf{H}^{-1}$. The resulting algorithm is GPTQ, which can efficiently quantize billion-parameter LLMs.
@@ 3 Methodology
Figure 1: (Left) The empirical distributions of inverse Hessian diagonals, computed on 262K tokens from the C4 dataset for the Llama-3-8B model, contain outliers that can cause high loss errors. (Right) Our proposed loss-error-aware non-uniform and affine grids better preserve the quantized precision of outliers, leading to more accurate quantized models.
@@ 3.1 Revisiting the Loss Error
To motivate our proposed approach, we first revisit the loss error $\epsilon_{i}$ in Equation 3, which approximates the (detrimental) increase in the network’s task loss, introduced by quantizing weight $w_{i}$. This error $\epsilon_{i}$ has been used as a heuristic in multiple previous works (obd; obs; woodfisher; obq) for choosing the next best weight $i$ to prune or quantize. It has been shown to be a highly informative metric for measuring the impact of quantization.
@@
Hence, we further examine the empirical distribution of $\frac{1}{\mathrm{diag}(\mathbf{H}^{-1})}$, which is proportional to $\bm{\epsilon}$, the loss error of an entire row. We obtain the empirical distributions on layers of Llama-3-8B (llama3) with 128 sequences of length 2048 tokens from the C4 dataset (c4), and compute the inverse Hessian as $\mathbf{H}^{-1}=(2\mathbf{X}\mathbf{X}^{\top})^{-1}$ where $\mathbf{X}$ is the layer input matrix. As shown in Figure 1, The majority of the inverse diagonals are concentrated in low-magnitude regions, with a few outliers having high magnitudes. Quantizing the weights corresponding to these outliers can lead to high loss errors if these weights are not well-aligned with the quantization grid points. Preserving the quantized precision of the weights corresponding to these inverse-diagonal outliers is especially important because the loss error increases quadratically with their quantization error (Equation 4). Iterative loss-error-based quantization approaches (OBQ, GPTQ, etc.) employ min-max affine quantization grid, which is suboptimal for preserving the quantized precision of the inverse-diagonal outliers, leading to high loss errors and model quality degradation. Our idea is to learn quantization grids that minimize the loss error $\epsilon$.